How much heat would be absorbed by 2 litres of an ideal gas at 5 atm pressure if it expands isothermally into an evacuated flask of 5 litres

Grade:12th pass

1 Answers

Siddhant
6 years ago
In the case of expansion of any gas, work is done by the system.Now we know that w=−pextΔV=−pext(Vf−Vi)where pext​ = external pressure of the system = 0(The given pressure of 10 atm in the question, is the pressure of the gas itself. But we need external pressure in the expression. The gas is expanding into vacuum which has no pressure. Thus external pressure = 0) Vf = 10 L Vi = 2 LThus work done by the system will be: w = -0×(10-2) = 0Thus no work is done.Also, heat absorbed (q) is a consequence of gaining temperature, but the system is working at constant temperature (isothermal condition).q = -wSo there will be no heat absorbed.